# Homogeneous braids are visually prime

### v1

@@ 6. Trefoil plumbings do not preserve primeness
We will now prove Proposition 1.4, which involves the construction of an open book $(\Sigma_{P},P)$ and an essential arc $a$ such that trefoil plumbing along $a$ yields an open book that has a fixed essential separating arc $\gamma$. Further modifications will allow us to see that we can arrange for the binding to be connected and for the original open book to have no fixed essential arcs.
@@
Consider the genus 2 surface with two boundary components. We denote it by $\Sigma_{P}$ and pick an explicit model: we view $\Sigma_{P}$ as the result of surgering one-holed tori on to an annulus at two spots given by two marked points; see left-hand side of Figure 12.
@@
Figure 12. Left: The surface $\Sigma_{P}$, depicted as an annulus with two discs (black) which indicate where two once punctured tori are surgered in, and three essential arcs $\alpha$ (dotted, blue), $\beta_{1}$ (blue), and $\beta_{2}$ (blue). Also an immersed circle $C$ (green) in the annulus that runs through both discs. Right: The images of $\alpha$, $\beta_{1}$, and $\beta_{2}$ under $P$, and the stabilization arc $a$ (red).
@@
We also pick three essential arcs $\alpha$, $\beta_{1}$, and $\beta_{2}$ in $\Sigma_{P}$. Those arcs union the cores of the handle of the plumbings will form the fixed arc $\gamma$.
@@
We define the monodromy $P$ in two steps. Firstly, we define it as a diffeomorphism of the annulus with two marked points. For this we pick $P(\alpha)$, $P(\beta_{1})$, and $P(\beta_{2})$ as depicted on the right-hand side of Figure 12. There exists a diffeomorphism mapping $\alpha$, $\beta_{1}$, and $\beta_{2}$ to $P(\alpha)$, $P(\beta_{1})$, and $P(\beta_{2})$ that fixes the boundary pointwise and the two marked points as a set, and this diffeomorphism, which we denote by $P$, is unique up to isotopy relative boundary. To see this, we note that $\alpha$, $\beta_{1}$, $\beta_{2}$ cut the annulus into a disc (in fact a $6$–gon) and discs with one marked point (one bigon and one punctured $4$–gon each with one marked point), and the same is true for $P(\alpha)$, $P(\beta_{1})$, $P(\beta_{2})$. [^1] As an alternative description of $P$, the reader may consider the immersed oriented circle $C$ with a marked point in the annulus depicted on the left-hand side of Figure 12, and take $P$ to be the half-push map that exchanges the two marked points, starting with the one marked on $C$ and pushing in the direction of the orientation. We will not use the latter description of $P$. Secondly, we choose an extension of $P$ to a self–diffeomorphism of $\Sigma_{P}$, which, by abuse of notation, we denote as $P\colon\Sigma_{P}\to\Sigma_{P}$. By construction, $P$ interchanges the two one-holed tori surgered to the annulus at the marked points.
@@
Next, we choose an essential arc $a$ as depicted in the right-hand side of Figure 12, along which we perform a positive stabilization yielding an open book $(\Sigma_{1},\phi_{1})$. By viewing $\Sigma_{P}$ as a subsurface of $\Sigma_{1}$, $\Sigma_{1}$ naturally features a copy of the arc $\alpha$ but furthermore features an arc $\beta$ given as the union of $\beta_{1}$ and $\beta_{2}$ with a core of the newly attached handle $\Sigma_{1}\setminus\mathrm{int}(\Sigma_{P})$; see the left of Figure 13. Extending $P$ to $\Sigma_{1}$ by the identity, the resulting monodromy $\phi_{1}$ can be understood as the composition of $P$ with a positive Dehn twist $\tau_{a^{\prime}}$ along the simple closed curve given as the union of $a$ with a core of the new handle. The result of applying $\phi_{1}$ to $\alpha$ and $\beta$ is depicted on the right in Figure 13.
@@
Figure 13. Left: the arcs $\alpha$ and $\beta$ on $\Sigma_{1}$, together with the curve $a^{\prime}$ (obtained by joining $a$ with the core of the handle). The new handle corresponding to the stabilization is bounded by the dashed lines. Right: the images $\phi_{1}(\alpha)=\tau_{a^{\prime}}(P(\alpha))$ and $\phi_{1}(\beta)=\tau_{a^{\prime}}(P(\beta))$; and the cocore of the handle $b$ along which we perform the second stabilization.
@@
We also consider the arc $b$ in $\Sigma_{1}$ as given in Figure 13, and let $(\Sigma,\phi)$ be the open book resulting from $(\Sigma_{1},\phi_{1})$ by positive stabilization along $b$. Since $b$ is a cocore of the handle $\Sigma_{1}\setminus\mathrm{int}(\Sigma_{P})$, the open book $(\Sigma,\phi)$ is equivalent to the trefoil plumbing of $(\Sigma_{P},P)$ along $a$. We next describe a fixed essential separating arc $\gamma$ of $(\Sigma,\phi)$.
@@
Thinking of $\Sigma_{1}$ as a subset of $\Sigma$, note that $\alpha$ and $\beta$ are chosen such that they each have one endpoint in $\partial\Sigma_{1}$ that lies in the interior of $\Sigma$. We define the arc $\gamma$ in $\Sigma$ to be the union of $\alpha$, $\beta$, and a core $h$ of the newly attached handle $\Sigma\setminus\mathrm{int}(\Sigma_{1})$, see the left hand side of Figure 14. One readily checks that $\gamma$ is an essential separating arc, and it is not hard to see that $\gamma$ is fixed. Indeed, note for example $\phi_{1}(\gamma)$, which equals the union of $\phi_{1}(\alpha)\cup\phi_{1}(\beta)\cup h$, is isotopic to the result of applying $\phi_{2}^{-1}$. Here $\phi_{1}\colon\Sigma\to\Sigma$ is the extension by the identity of the monodromy of $(\Sigma_{1},\phi_{1})$ and $\phi_{2}\colon\Sigma\to\Sigma$ is the extension by the identity of the positive Dehn twist that is the monodromy of the Hopf band that was plumbed in the stabilization along $b$. A representative of $\phi_{1}(\gamma)=\phi_{2}^{-1}(\gamma)$, chosen so that it differs from $\gamma$ only in the plumbing region, can be seen on the right-hand side of Figure 14.
@@
It remains to discuss that the open book $(\Sigma_{P},P)$ has no fixed essential arcs, and to establish that we can arrange for the open book to have connected binding.
@@
For the latter, assuming $(\Sigma_{P},P)$ has no essential fixed arcs, we can consider the open book $(\Sigma_{0},\phi_{0})$ resulting from iteratively plumbing three Hopf bands to $(\Sigma_{P},P)$ along an arc $c$ connecting two boundary components such that $c$ is disjoint from $a$, $P(\alpha)$, $P(\beta_{1})$, $P(\beta_{2})$. More precisely, consider the tree $T$ of open books with vertices $v_{1}=(\Sigma_{P},P)$, $v_{2}$, $v_{3}$, and $v_{4}$, where $v_{2}$ and $v_{4}$ are open books of positive Hopf bands, while $v_{3}$ is the open book of a negative Hopf band, and take the edge set to be $\{\{v_{1},v_{2}\},\{v_{2},v_{3}\},\{v_{3},v_{4}\}\}$. Further take the labelings $P_{1}$, $P_{2}$, and $P_{3}$ such that $P_{1}$ is the plumbing region in $\Sigma_{P}$ given as a $4$–gon that is a neighborhood of $c$ (all other plumbing regions are uniquely determined up to isotopy, since in annuli there is only one essential $4$–gon plumbing region up to isotopy). We claim that the open book $(\Sigma_{0},\phi_{0})$ (well-defined up to equivalence) associated with $T$ has no fixed essential arc. To see this we note that, if $T$ had one less vertex, namely $v_{4}$, and one less edge, namely $\{v_{3},v_{4}\}$, then the corresponding open book $(\Sigma_{0}^{\prime},\phi_{0}^{\prime})$ is an essential figure-eight plumbing, which has no fixed essential arc by Proposition 1.5. Given that $T$ has the extra vertex and edge, the associated open book is in fact the result of positive stabilization of $(\Sigma_{0}^{\prime},\phi_{0}^{\prime})$. Now, the plumbing arc for the stabilization veers to the left in $(\Sigma_{0}^{\prime},\phi_{0}^{\prime})$, $(\Sigma_{0}^{\prime},\phi_{0}^{\prime})$ has no fixed essential arcs, and the stabilization is positive (meaning we are summing a strictly right-veering open book to $(\Sigma_{0}^{\prime},\phi_{0}^{\prime})$). Thus, 1.3 implies that the open book $(\Sigma_{0},\phi_{0})$ has no fixed essential arcs. By construction $\Sigma_{0}$ has connected boundary; hence $(\Sigma_{0},\phi_{0})$ is an open book with connected binding as desired.
@@
Finally, we slightly modify $(\Sigma_{P},P)$ to assure it does not contain fixed essential arcs. First, we argue that we may assume that all essential fixed arcs are disjoint from the interior of $\alpha$, $P(\alpha)$ $P(\beta_{1})$, and $P(\beta_{2})$. To see this, we simply modify $(\Sigma_{P},\phi_{P})$ by performing figure-eight plumbings along (parallel copies of) the essential arcs $P(\alpha)$, $P(\beta_{1})$, and $P(\beta_{2})$. The result is a new open book $(\Sigma_{0},\phi_{0})$ with the property that all essential fixed arcs can be isotoped to miss the plumbing arcs by Remark 4.2. In particular, if $(\Sigma_{0},\phi_{0})$ has an essential fixed arc, then $(\Sigma_{P},\phi_{P})$ has an essential fixed arc disjoint from $P(\beta_{1})$, $P(\beta_{2})$, and $P(\alpha)$. Furthermore, note that $\phi_{0}$ and $P$ map $\alpha$, $\beta_{1}$, and $\beta_{2}$ to the same curves since we chose figure-eight plumbing arcs that are disjoint from $P(\alpha)$, $P(\beta_{1})$, and $P(\beta_{2})$. Hence, the construction of the fixed arc $\gamma$ goes through for $(\Sigma_{0},\phi_{0})$ in the exact same way as for $(\Sigma_{P},\phi_{P})$.
@@
With this we have reduced to the problem of showing that $(\Sigma_{P},\phi_{P})$ has no fixed essential arc that is disjoint from $P(\beta_{1})$, $P(\beta_{2})$ and $P(\alpha)$. We observed earlier that $\Sigma_{P}\setminus(P(\beta_{1})\cup P(\beta_{2})\cup P(\alpha))$ has three connected components. We discuss that none of these contain fixed essential arcs. One component is a $6$–gon bounded by segments in $\partial\Sigma_{P}$ and all three arcs $P(\beta_{1})$, $P(\beta_{2})$, $P(\alpha)$. Every arc in this $6$–gon that is not boundary parallel in $\Sigma_{P}$ is parallel to one of $P(\beta_{1})$, $P(\beta_{2})$, $P(\alpha)$, none of which are fixed. A second component is a bigon with a marked point, where the marked point is replaced by a one-holed torus, that is bounded by a segment in $\partial\Sigma_{P}$ and $P(\beta_{1})$. A fixed essential arc $\gamma$ must necessarily enter the one-holed torus, as otherwise it is parallel to the boundary (and hence not essential) or to $P(\beta_{1})$, and the latter is not fixed. Here, entering the one-holed torus is formalized as having non-trivial intersection with the simple closed curve encircling the marked point. However, such a $\gamma$ is disjoint from the curve that encircles the other marked point—respectively one-holed torus—, but $P(\gamma)$ intersects said curve since $P$ exchanges the two curves that encircle the marked points. In particular, $\gamma$ and $P(\gamma)$ are not isotopic. The third component is a $4$–gon with a marked point, where the marked point is replaced by a one-holed torus, that is bounded by segments in $\partial\Sigma_{P}$, $P(\beta_{2})$ and $P(\alpha)$. A similar argument as above shows that any fixed essential $\gamma$ must enter the one-holed torus, but does, of course not enter the other one-holed torus, hence $\gamma$ cannot be fixed as $P$ exchanges the two one-holed tori. This concludes the proof that we may choose the open book so that it does not contain fixed essential arcs. ∎

### v2

@@ 6. Trefoil plumbings do not preserve primeness
We will now prove Proposition 1.4, which involves the construction of an open book $(\Sigma_{\rho},\rho)$ and an essential arc $a$ such that trefoil plumbing along $a$ yields an open book that has a fixed essential separating arc $\gamma$. Further modifications will allow us to see that we can arrange for the binding to be connected and for the original open book to have no fixed essential arcs.
@@
Consider the genus 2 surface with two boundary components. We denote it by $\Sigma_{\rho}$ and pick an explicit model: we view $\Sigma_{\rho}$ as the result of surgering one-holed tori onto an annulus at two spots given by two marked points; see left-hand side of Figure 12.
@@
Figure 12. Left: The surface $\Sigma_{\rho}$, depicted as an annulus with two discs (black) which indicate where two once punctured tori are surgered in, and three essential arcs $\alpha$ (dotted, blue), $\beta_{1}$ (blue), and $\beta_{2}$ (blue). Also an immersed circle $C$ (green) in the annulus that runs through both discs. Right: The images of $\alpha$, $\beta_{1}$, and $\beta_{2}$ under $\rho$, and the stabilization arc $a$ (red).
@@
We also pick three essential arcs $\alpha$, $\beta_{1}$, and $\beta_{2}$ in $\Sigma_{\rho}$. We will use those arcs together with the cores of the handle of the plumbings to obtain the fixed arc $\gamma$.
@@
We define the monodromy $\rho$ in two steps. Firstly, we define it as a diffeomorphism of the annulus with two marked points. For this we pick $\rho(\alpha)$, $\rho(\beta_{1})$, and $\rho(\beta_{2})$ as depicted on the right-hand side of Figure 12. There exists a diffeomorphism mapping $\alpha$, $\beta_{1}$, and $\beta_{2}$ to $\rho(\alpha)$, $\rho(\beta_{1})$, and $\rho(\beta_{2})$ that fixes the boundary pointwise and the two marked points as a set, and this diffeomorphism, which we denote by $\rho$, is unique up to isotopy relative boundary. To see this, we note that $\alpha$, $\beta_{1}$, $\beta_{2}$ cut the annulus into a disc (in fact a 66–gon) and discs with one marked point (one bigon and one punctured 44–gon each with one marked point), and the same is true for $\rho(\alpha)$, $\rho(\beta_{1})$, $\rho(\beta_{2})$. [^2] As an alternative description of $\rho$, the reader may consider the immersed oriented circle $C$ with a marked point in the annulus depicted on the left-hand side of Figure 12, and take $\rho$ to be the half-push map that exchanges the two marked points, starting with the one marked on $C$ and pushing in the direction of the orientation. We will not use the latter description of $\rho$. Secondly, we choose an extension of $\rho$ to a self–diffeomorphism of $\Sigma_{\rho}$, which, by abuse of notation, we denote as $\rho\colon\Sigma_{\rho}\to\Sigma_{\rho}$. By construction, $\rho$ interchanges the two one-holed tori surgered to the annulus at the marked points.
@@
Next, we choose an essential arc $a$ as depicted in the right-hand side of Figure 12, along which we perform a positive stabilization yielding an open book $(\Sigma_{1},\phi_{1})$. By viewing $\Sigma_{\rho}$ as a subsurface of $\Sigma_{1}$, $\Sigma_{1}$ naturally features a copy of the arc $\alpha$ but furthermore features an arc $\beta$ given as the union of $\beta_{1}$ and $\beta_{2}$ with a core of the newly attached handle $\Sigma_{1}\setminus\mathrm{int}(\Sigma_{\rho})$; see the left of Figure 13. Extending $\rho$ to $\Sigma_{1}$ by the identity, the resulting monodromy $\phi_{1}$ can be understood as the composition of $\rho$ with a positive Dehn twist $\tau_{a^{\prime}}$ along the simple closed curve given as the union of $a$ with a core of the new handle. The result of applying $\phi_{1}$ to $\alpha$ and $\beta$ is depicted on the right in Figure 13.
@@
Figure 13. Left: the arcs $\alpha$ and $\beta$ on $\Sigma_{1}$, together with the curve $a^{\prime}$ (obtained by joining $a$ with the core of the handle). The new handle corresponding to the stabilization is bounded by the dashed lines. Right: the images $\phi_{1}(\alpha)=\tau_{a^{\prime}}(\rho(\alpha))$ and $\phi_{1}(\beta)=\tau_{a^{\prime}}(\rho(\beta))$; and the cocore of the handle $b$ along which we perform the second stabilization.
@@
We also consider the arc $b$ in $\Sigma_{1}$ as given in Figure 13, and let $(\Sigma,\phi)$ be the open book resulting from $(\Sigma_{1},\phi_{1})$ by positive stabilization along $b$. Since $b$ is a cocore of the handle $\Sigma_{1}\setminus\mathrm{int}(\Sigma_{\rho})$, the open book $(\Sigma,\phi)$ is equivalent to the trefoil plumbing of $(\Sigma_{\rho},\rho)$ along $a$. We next describe a fixed essential separating arc $\gamma$ of $(\Sigma,\phi)$.
@@
Note that $\alpha$ and $\beta$ are chosen such that they each have one endpoint in a corner of the plumbing region. Thinking of $\Sigma_{1}$ as a subset of $\Sigma$, by a small isotopy we can assume that the arcs have that endpoint in a part of $\partial\Sigma_{1}$ that lies in the interior of $\Sigma$. We define the arc $\gamma$ in $\Sigma$ to be the union of $\alpha$, $\beta$, and a core $h$ of the newly attached handle $\Sigma\setminus\mathrm{int}(\Sigma_{1})$, see the left hand side of Figure 14. One readily checks that $\gamma$ is an essential separating arc, and it is not hard to see that $\gamma$ is fixed. Indeed, note for example $\phi_{1}(\gamma)$, which equals the union of $\phi_{1}(\alpha)\cup\phi_{1}(\beta)\cup h$, is isotopic to the result of applying $\phi_{2}^{-1}$. Here $\phi_{1}\colon\Sigma\to\Sigma$ is the extension by the identity of the monodromy of $(\Sigma_{1},\phi_{1})$ and $\phi_{2}\colon\Sigma\to\Sigma$ is the extension by the identity of the positive Dehn twist that is the monodromy of the Hopf band that was plumbed in the stabilization along $b$. A representative of $\phi_{1}(\gamma)=\phi_{2}^{-1}(\gamma)$, chosen so that it differs from $\gamma$ only in the plumbing region, can be seen on the right-hand side of Figure 14.
@@
It remains to discuss that the open book $(\Sigma_{\rho},\rho)$ has no fixed essential arcs, and to establish that we can arrange for the open book to have connected binding.
@@
For the latter, assuming $(\Sigma_{\rho},\rho)$ has no essential fixed arcs, we can consider the open book $(\Sigma_{0},\phi_{0})$ resulting from iteratively plumbing three Hopf bands to $(\Sigma_{\rho},\rho)$ along an arc $c$ connecting two boundary components such that $c$ is disjoint from $a$, $\rho(\alpha)$, $\rho(\beta_{1})$, $\rho(\beta_{2})$. More precisely, consider the tree $T$ of open books with vertices $v_{1}=(\Sigma_{\rho},\rho)$, $v_{2}$, $v_{3}$, and $v_{4}$, where $v_{2}$ and $v_{4}$ are open books of positive Hopf bands, while $v_{3}$ is the open book of a negative Hopf band, and take the edge set to be $\{\{v_{1},v_{2}\},\{v_{2},v_{3}\},\{v_{3},v_{4}\}\}$. Further take the labelings $P_{1}$, $P_{2}$, and $P_{3}$ such that $P_{1}$ is the plumbing region in $\Sigma_{\rho}$ given as a 44–gon that is a neighborhood of $c$ (all other plumbing regions are uniquely determined up to isotopy, since in annuli there is only one essential 44–gon plumbing region up to isotopy). We claim that the open book $(\Sigma_{0},\phi_{0})$ (well-defined up to equivalence) associated with $T$ has no fixed essential arc. To see this we note that, if $T$ had one less vertex, namely $v_{4}$, and one less edge, namely $\{v_{3},v_{4}\}$, then the corresponding open book $(\Sigma_{0}^{\prime},\phi_{0}^{\prime})$ is an essential figure-eight plumbing, which has no fixed essential arc by Proposition 1.5. Given that $T$ has the extra vertex and edge, the associated open book is in fact the result of positive stabilization of $(\Sigma_{0}^{\prime},\phi_{0}^{\prime})$. Now, the plumbing arc for the stabilization veers to the left in $(\Sigma_{0}^{\prime},\phi_{0}^{\prime})$, $(\Sigma_{0}^{\prime},\phi_{0}^{\prime})$ has no fixed essential arcs, and the stabilization is positive (meaning we are summing a strictly right-veering open book to $(\Sigma_{0}^{\prime},\phi_{0}^{\prime})$). Thus, 1.3 implies that the open book $(\Sigma_{0},\phi_{0})$ has no fixed essential arcs. By construction $\Sigma_{0}$ has connected boundary; hence $(\Sigma_{0},\phi_{0})$ is an open book with connected binding as desired.
@@
Finally, we slightly modify $(\Sigma_{\rho},\rho)$ to assure it does not contain fixed essential arcs. First, we argue that we may assume that all essential fixed arcs are disjoint from the interior of $\alpha$, $\rho(\alpha)$ $\rho(\beta_{1})$, and $\rho(\beta_{2})$. To see this, we simply modify $(\Sigma_{\rho},\rho)$ by performing figure-eight plumbings along (parallel copies of) the essential arcs $\rho(\alpha)$, $\rho(\beta_{1})$, and $\rho(\beta_{2})$. The result is a new open book $(\Sigma_{0},\phi_{0})$ with the property that all essential fixed arcs can be isotoped to miss the plumbing arcs by Remark 4.2. In particular, if $(\Sigma_{0},\phi_{0})$ has an essential fixed arc, then $(\Sigma{\rho}P,\rho)$ has an essential fixed arc disjoint from $\rho(\beta_{1})$, $\rho(\beta_{2})$, and $\rho(\alpha)$. Furthermore, note that $\phi_{0}$ and $\rho$ map $\alpha$, $\beta_{1}$, and $\beta_{2}$ to the same curves since we chose figure-eight plumbing arcs that are disjoint from $\rho(\alpha)$, $\rho(\beta_{1})$, and $\rho(\beta_{2})$. Hence, the construction of the fixed arc $\gamma$ goes through for $(\Sigma_{0},\phi_{0})$ in the exact same way as for $(\Sigma_{\rho},\rho)$.
@@
With this we have reduced to the problem of showing that $(\Sigma_{\rho},\rho)$ has no fixed essential arc that is disjoint from $\rho(\beta_{1})$, $\rho(\beta_{2})$ and $\rho(\alpha)$. We observed earlier that $\Sigma_{\rho}\setminus(\rho(\beta_{1})\cup\rho(\beta_{2})\cup\rho(\alpha))$ has three connected components. We discuss that none of these contain fixed essential arcs. One component is a 66–gon bounded by segments in $\partial\Sigma_{\rho}$ and all three arcs $\rho(\beta_{1})$, $\rho(\beta_{2})$, $\rho(\alpha)$. Every arc in this 66–gon that is not boundary parallel in $\Sigma_{\rho}$ is parallel to one of $\rho(\beta_{1})$, $\rho(\beta_{2})$, $\rho(\alpha)$, none of which are fixed. A second component is a bigon with a marked point, where the marked point is replaced by a one-holed torus, that is bounded by a segment in $\partial\Sigma_{\rho}$ and $\rho(\beta_{1})$. A fixed essential arc $\gamma$ must necessarily enter the one-holed torus, as otherwise it is parallel to the boundary (and hence not essential) or to $\rho(\beta_{1})$, and the latter is not fixed. Here, entering the one-holed torus is formalized as having non-trivial intersection with the simple closed curve encircling the marked point. However, such a $\gamma$ is disjoint from the curve that encircles the other marked point—respectively one-holed torus—, but $\rho(\gamma)$ intersects said curve since $\rho$ exchanges the two curves that encircle the marked points. In particular, $\gamma$ and $\rho(\gamma)$ are not isotopic. The third component is a 44–gon with a marked point, where the marked point is replaced by a one-holed torus, that is bounded by segments in $\partial\Sigma_{\rho}$, $\rho(\beta_{2})$ and $\rho(\alpha)$. A similar argument as above shows that any fixed essential $\gamma$ must enter the one-holed torus, but does, of course not enter the other one-holed torus, hence $\gamma$ cannot be fixed as $\rho$ exchanges the two one-holed tori. This concludes the proof that we may choose the open book so that it does not contain fixed essential arcs. ∎